### v1

@@ 4 The Merge Phase
Our construction closely follows that of [racke2014computing]. For the readers familiar with [racke2014computing], we mention the main differences:
@@
    The “first partition” step (Theorem 4.1) is translated to the cut-sparsifier case (which improves the quality of the cut-sparsifier by a factor of $\log n$), and its proof is significantly simplified, as the “heavy algorithmic tools” of a balanced cut or expander theorem [ADK25] and a fair cuts computation [LNPSsoda23] are used in a black box manner.
@@
    The proof of the main lemma used in the second partition step (Lemma 4.6) is improved and simplified by using the results of [LNPSsoda23].
As explained in Section 2, the hierarchical decomposition is constructed by recursively applying the “merge phase” partitioning procedure to input clusters $S\subseteq V$. Each input cluster $S\subseteq V$ is partitioned in two levels. At the first level we define the direct children of $S$ to be disjoint clusters $L$ and $R$ such that $L\cup R=S$, and at the second level the children of $L$ and of $R$, respectively, are defined to be disjoint clusters $L_{1},\ldots,L_{\ell}$ and $R_{1},\ldots,R_{r}$ (satisfying $L_{1}\cup\cdots\cup L_{\ell}=L$ and $R_{1}\cup\cdots\cup R_{r}=R$). The size of the clusters $L_{i}$ and $R_{i}$ is bounded as $|L_{i}|\leq\frac{2}{3}|S|$ and $|R_{i}|\leq\frac{2}{3}|S|$.
@@
The intuition behind the two-level partition is as follows. As discussed in Section 2, we would want to partition $S$ into sub-clusters $\{S_{1},\ldots,S_{t}\}$ of size $|S_{i}|\leq c\cdot|S|,c<1$, so that the inter-cluster edges $F$ and the boundary edges $B$ of $S$, satisfy that $F\cup B$ expands. However, this may not be possible for a cluster $S$ if $B$ does not expand in $G[S]$. Consequently, the first partition into $Z_{1},\ldots,Z_{z}$ ignores $B$ and focuses on reducing the size of the clusters, while ensuring that $F$, the set of inter-cluster edges, expands. In Section 5, we describe how to move the demands that need to leave each cluster $S\in V_{T}$, to the cluster’s boundary edges. Specifically, we wish to move demands that reside on $F$ (i.e., boundary edges of sub-clusters) to $B$. Therefore, we use a fair-cut computation [LNPSsoda23] to find an approximate min-cut, $Y\subseteq E$, that blocks the flow problem of routing from $F$ to $B$. We then add this cut to the tree (i.e., find a partition $L\cup R$ induced by $Y$), thereby forcing the given demand matrix to respect this cut. This allows us to bound the amount of mass that has to cross the cut $Y$. See Section 5 for details on the usage of this partition.
@@ 4.1 Partition of $G[S]$
Theorem 4.1 (Improved version of [racke2014computing, Theorem 3.1]).
@@
    with high probability, the set of split vertices corresponding to the inter-cluster edges
@@
    $\alpha$-expands in $G[S]^{\prime}$ for $\alpha=\Omega\!\left(\frac{1}{\log n}\right)$ (recall the definition of the subdivision graph in Definition 3.1).
@@
Note that Theorem 4.1 has $\alpha=\Omega\!\left(\frac{1}{\log n}\right)$, instead of $\alpha=\Omega\!\left(\frac{1}{\log^{2}n}\right)$ in [racke2014computing, Theorem 3.1], which leads (as shown below) to the improvement by a factor of $\log n$ in the quality of the tree cut-sparsifier. The partitioning algorithm in Theorem 4.1 is independent of the edges outside $G[S]$. Therefore, we can restrict our attention to the induced subgraph $G[S]$: in the rest of this section, $G=(V,E)$ refers to $G[S]$, and $n,m$ denote the number of vertices and edges, respectively, in $G[S]$. Following the notation of [racke2014computing], we say that a subset $F\subseteq E$ of edges induces a balanced clustering in $G$, if deleting $F$ from $G$ leaves every connected component with at most $\frac{2}{3}|V|$ vertices.
To establish Theorem 4.1, we first show the following lemma, which improves and unifies Lemmas 3.1 and 3.2 in [racke2014computing].
Lemma 4.2 (Improved version of [racke2014computing, Lemmas 3.1 and 3.2]).
@@
    an edge set $\tilde{F}=\tilde{A}\cup\tilde{C}$, such that $\tilde{F}$ induces a balanced clustering, and with high probability $X_{\tilde{A}}$ $\Omega(1/\log n)$-expands in $G^{\prime}$. Additionally there exists a flow in $G^{\prime}$ from nodes in $X_{\tilde{C}}$ to $X_{\tilde{A}}$ that can be routed with congestion $O(1)$ and:
@@
We begin by showing how Lemma 4.2 implies Theorem 4.1.
@@
Once Lemma 4.2 terminates with Case (2), $X_{\tilde{A}}$ $\Omega(1/\log n)$-expands in $G^{\prime}$, and we can route in $G^{\prime}$ with congestion $O(1)$ a flow such that each node in $X_{\tilde{C}}$ sends $1$ unit of flow and each node in $X_{\tilde{A}}$ receives at most $1$ unit of flow. By Corollary 4.4 below, this implies that edges in $\tilde{F}=\tilde{A}\cup\tilde{C}$ satisfy that $X_{\tilde{F}}$ $\Omega(1/\log n)$-expands in $G^{\prime}$. Additionally, $\tilde{F}$ induces a balanced clustering whose connected components will be $\{Z_{1},\ldots,Z_{z}\}$. [^6] After computing $\tilde{F}$ and the resulting partition $\{Z_{1},\ldots,Z_{z}\}$, we re-define $F$ to be the new set of inter-cluster edges $F\mathrel{\mathop{:}}=\{(u,v)\mid u\in Z_{i},v\in Z_{j},(u,v)\in E,i\neq j\}$. This set satisfies $F\subseteq\tilde{F}$ and thus $\Omega(1/\log n)$-expands in $G[S]^{\prime}$ (see Fact 3.8). This proves Theorem 4.1. ∎
Next, we prove that $X_{\tilde{F}}=X_{\tilde{A}}\cup X_{\tilde{C}}$ $\Omega(1/\log n)$-expands in $G^{\prime}$. Intuitively (see the discussion proceeding Definition 3.5), we wish to show that $G^{\prime}$ respects the all-to-all multi-commodity flow problem between the nodes of $X_{\tilde{F}}$ (see Definition 2.2). Using the fact that $X_{\tilde{A}}$ expands in $G^{\prime}$ (thus $G^{\prime}$ respects the all-to-all flow problem on $X_{\tilde{A}}$) and that there is a low congestion flow $f$ from $X_{\tilde{C}}$ to $X_{\tilde{A}}$ (in particular $G^{\prime}$ respects the flow problem that moves mass from $X_{\tilde{C}}$ to $X_{\tilde{A}}$ according to $f$), we prove that $G^{\prime}$ respects the flow problem (i.e., demand matrix) corresponding to the following transition of mass:
@@
    First, “move” the demands in $X_{\tilde{C}}$ to $X_{\tilde{A}}$ according to $f$.
@@
    Then, spread the demands in $X_{\tilde{A}}$ uniformly (using the fact that $X_{\tilde{A}}$ expands in $G^{\prime}$).
@@
    Finally, “move” part of the (uniform) demands in $X_{\tilde{A}}$ back to $X_{\tilde{C}}$ according $f$ (in reverse).
We show a slightly more general result:
Lemma 4.3.
Let $G$ be a graph, and let $A,C\subseteq V$. Assume that AA $\phi$-expands in $G$, and additionally that there exists a flow routable with congestion $c$, such that each vertex $v\in C$ sends $1$ unit of flow, and each vertex $u\in A$ receives at most $a$ units. Then $A\cup C$ $\frac{\phi}{2\cdot(a+1+c\phi)}$-expands in $G$.
Proof.
@@
Consider the all-to-all flow problem (according to Definition 2.2) $Q_{A\cup C}$ on $A\cup C$ in $G$. According to Remark 3.6, we should show that $G$ respects $Q_{A\cup C}$. Let $P_{1}\mathrel{\mathop{:}}=P_{A\cup C}$ be the corresponding valid demand state (see Remark 3.15). Note that for each $u\in A\cup C$, $\left\|{P_{1}}(u)\right\|_{1}=1-\frac{1}{|A\cup C|}\leq 1$. Now, we use the given flow from $C$ to $A$, denoted $Q$, to transport the demand state vectors of the vertices in $C\setminus A$ to vertices in $A$. We scale the flow paths of $Q$ (i.e., the rows of its demand matrix) such that each $u\in C\setminus A$ sends $1-\frac{1}{|A\cup C|}$ units and the rest send 0. Let $Q^{\prime}$ be the scaled flow, and consider the updated demand state $P_{2}\mathrel{\mathop{:}}=P_{1}^{\uparrow Q^{\prime}}$ (recall Definition 3.18). Similar to Remark 3.20, $P_{2}$ is now supported only on $A$. Additionally, the load on each $u\in A$ is at most $a+1$ units. Because $Q$ can be routed with congestion $c$ and $Q^{\prime}$ is a scaled down variant of $Q$, we get that in particular (see Fact 2.7) $G$ $\frac{1}{c}$-respects $Q^{\prime}$, so by Fact 3.19, for any cut $(B,W=V\setminus B)$ we have $\operatorname{dem}_{(P_{1}-P_{2})}(B,W)\leq c\cdot\operatorname{cap}_{G}(B,W)$.
@@
Next, since AA $\phi$-expands in $G$, Remark 3.6 implies that $G$ $\phi$-respects the all-to-all flow on $A$, denoted $Q_{A}$. Scaling this flow problem by at most $a+1$, we can get a flow problem which routes $\left\|{P_{2}}(u)\right\|_{1}$ units out of each $u\in A$, and spreads them uniformly on $A$. This scaled problem is $\frac{\phi}{a+1}$-respected by $G$. Updating $P_{2}$ using this scaled flow gives $P_{3}$, which by Remark 3.20, has the same demand vectors ${P_{3}}(u)$ for all $u\in A$. Because $P_{1}$ is a valid demand state, so are $P_{2}$ and $P_{3}$ (see Fact 3.19). This means that $P_{3}\equiv 0$.
@@
Similar to before, we get that for any cut $(B,W)$ we have $\operatorname{dem}_{(P_{2}-P_{3})}(B,W)\leq\frac{a+1}{\phi}\cdot\operatorname{cap}_{G}(B,W)$. Therefore, using Fact 3.17, for any cut $(B,W)$, we have
$\operatorname{dem}_{Q_{A\cup C}}(B,W)$ $=\operatorname{dem}_{P_{1}}(B,W)\leq\operatorname{dem}_{(P_{1}-P_{2})}(B,W)+\operatorname{dem}_{(P_{2}-P_{3})}(B,W)$
$\leq\left(c+\frac{a+1}{\phi}\right)\cdot\operatorname{cap}_{G}(B,W)=\frac{a+1+c\phi}{\phi}\cdot\operatorname{cap}_{G}(B,W),$
@@
As a direct corollary of Lemmas 4.2 and 4.3, we get:
Corollary 4.4.
The set $X_{\tilde{F}}$ $\Omega(1/\log n)$-expands in $G^{\prime}$.
@@
Figure 2: Partitioning of cluster $S$. Figures (a)-(c) represent $G$, figures (d)-(f) represent the split graph. Black ovals depict vertices and squares depict split nodes. Figures $(b)$ and $(e)$ correspond to the first partition. Blue edges (and blue squares) depict inter-cluster edges of the first partition (that is, $F$ and $X_{F}$, respectively). Figure $(f)$ depicts the second partition (which is computed in $G^{\prime}[S^{\prime}]$). The core $L$ is enveloped in gold. Split nodes in $X_{Y}$ are filled with gold. Split nodes in $X_{F\cap Y}$ are colored in blue and gold.
To prove Lemma 4.2 we need the following claim. It is a slightly more general rephrasing of Claim 1 of [racke2014computing]. We provide its proof for completeness.
Claim 4.5 ([racke2014computing, Claim 1]).
@@
Given the subset $F\subseteq E$ that induces a balanced clustering, we use Corollary 3.11 on $G^{\prime}$ with $\phi\mathrel{\mathop{:}}=\Theta\left(\frac{1}{\log n}\right)$ and $\mu\mathrel{\mathop{:}}=\mu_{X_{F}}$. It terminates with one of three cases:
@@
    for some $c_{1}>0$. Let $C^{\prime}\subseteq E^{\prime}$ be the set of edges that cross the cut $(A^{\prime},\bar{A^{\prime}})$ in $G^{\prime}$. Let $C\subseteq E$ be the corresponding set in $G$: $C\mathrel{\mathop{:}}=\{e\in E\mid\exists c\in C^{\prime}:x_{e}\in c\}$. We have $|C|\leq|C^{\prime}|$. Inequality (1) is equivalent to $|C^{\prime}|\leq\frac{1}{2}\min(|X_{F}\cap A^{\prime}|,|X_{F}\setminus A^{\prime}|)$. Let $A\subseteq E$ be the set of edges corresponding to $A^{\prime}$: $A\mathrel{\mathop{:}}=\{e\in E\mid x_{e}\in A^{\prime}\}$. We have $|F\cap A|=|X_{F}\cap A^{\prime}|$ and $|F\setminus A|=|X_{F}\setminus A^{\prime}|$. Therefore,
@@
    We find a cut $(A^{\prime},\bar{A^{\prime}})$ in $G^{\prime}$ with $0<|X_{F}\setminus A^{\prime}|\leq\frac{|F|}{\log n}$, $\Phi_{G^{\prime}}^{X_{F}}(A^{\prime},\bar{A^{\prime}})=O(\phi\log n)$, and with high probability, $A^{\prime}$ $\Omega(\phi)$-expands with respect to $X_{F}$ in $G$. As $\phi=\Theta\left(\frac{1}{\log n}\right)$, denote by $c_{2}>0$ the constant such that $\Phi_{G^{\prime}}^{X_{F}}(A^{\prime},\bar{A^{\prime}})\leq c_{2}$. Like in the previous case, denote by $C^{\prime}\subseteq E^{\prime}$ the set of edges that cross the cut $(A^{\prime},\bar{A^{\prime}})$ in $G^{\prime}$. Let $C\subseteq E$ be the corresponding set in $G$: $C\mathrel{\mathop{:}}=\{e\in E\mid\exists c\in C^{\prime}:x_{e}\in c\}$. We have $|C|\leq|C^{\prime}|$. Because the cut is sparse with respect to $\mu_{X_{F}}$, $|C^{\prime}|\leq c_{2}\cdot|X_{F}\setminus A^{\prime}|$.
    From remark 3.10, we get that there exists a flow in $G^{\prime}$, routable with congestion $O(1)$, such that each split-node in $X_{C}$ sends $1$ unit of flow and each node $v\in A^{\prime}$ receives at most $O(\phi\cdot\mu(v))$ units. Recall that $\phi=\Theta\left(\frac{1}{\log n}\right)$, so for large enough $n$, we may assume that each node receives at most $\mu(v)$ units of flow. In particular, only split nodes $v\in X_{F}$ can receive flow, and each receives at most one unit.
@@
      Assume $\tilde{R}\cup C$ induces a balanced clustering. In this case, we will set $\tilde{F}\mathrel{\mathop{:}}=\tilde{R}\cup C$, and claim that we get Case (1) of the lemma. Indeed, we have $|\tilde{R}\cup C|\leq|\tilde{R}|+|C|=|F\setminus A|+|C|\leq(c_{2}+1)\cdot|X_{F}\setminus A^{\prime}|\leq(c_{2}+1)\cdot\frac{|F|}{\log n}<|F|\cdot(1-\frac{1}{\log n})$ for large enough $n$.

### v2

@@ 4 The Merge Phase
Our construction closely follows that of [27]. For the readers familiar with [27], we mention the main differences:
@@
    The “first partition” step (Theorem 4.1) is translated to the cut-sparsifier case (which improves the quality of the cut-sparsifier by a factor of $\log n$), and its proof is significantly simplified, as the “heavy algorithmic tools” of a balanced cut or expander theorem [2] and a fair cuts computation [21] are used in a black box manner.
@@
    The proof of the main lemma used in the second partition step (Lemma 4.5) is improved and simplified by using the results of [21].
As explained in Section 2, the hierarchical decomposition is constructed by recursively applying the “merge phase” partitioning procedure to input clusters $S\subseteq V$. Each input cluster $S\subseteq V$ is partitioned in two levels. At the first level, we define the direct children of $S$ to be disjoint clusters $L$ and RR [^6] In may happen that $R=\emptyset$. In this case, no node corresponding to $R$ is added to the hierarchical tree $T$. Similarly, we do not add nodes corresponding to the empty set to the hierarchy. such that $L\cup R=S$, and at the second level the children of $L$ and of $R$, respectively, are defined to be disjoint clusters $L_{1},\ldots,L_{\ell}$ and $R_{1},\ldots,R_{r}$ (satisfying $L_{1}\cup\cdots\cup L_{\ell}=L$ and $R_{1}\cup\cdots\cup R_{r}=R$). The size of the clusters $L_{i}$ and $R_{i}$ is bounded as $|L_{i}|\leq\frac{2}{3}|S|$ and $|R_{i}|\leq\frac{2}{3}|S|$.
@@
The intuition behind the two-level partition is as follows. As discussed in Section 2, we want to partition $S$ into sub-clusters $\{S_{1},\ldots,S_{t}\}$ of size $|S_{i}|\leq c\cdot|S|,c<1$, so that the inter-cluster edges $F$ and the boundary edges $B$ of $S$, satisfy that $F\cup B$ expands. However, this may not be possible for a cluster $S$ if $B$ does not expand in $G[S]$. Consequently, the first partition into $Z_{1},\ldots,Z_{z}$ ignores $B$ and focuses on reducing the size of the clusters, while ensuring that $F$, the set of inter-cluster edges, expands. In Section 5, we describe how to move the demands that need to leave each cluster $S\in V_{T}$, to the cluster’s boundary edges. Specifically, we wish to move demands that reside on $F$ (i.e., boundary edges of sub-clusters) to $B$. Therefore, we use a fair-cut computation [21] to find an approximate min-cut, $Y\subseteq E$, that blocks the flow problem of routing from $F$ to $B$. We then add this cut to the tree (i.e., find a partition $L\cup R$ induced by $Y$), thereby forcing the given demand matrix to respect this cut. This allows us to bound the amount of mass that has to cross the cut $Y$. See Section 5 for details on the usage of this partition.
@@ 4.1 Partition of $G[S]$
Theorem 4.1 (Improved version of [27, Theorem 3.1]).
@@
    with high probability, in the subdivision graph $G[S]^{\prime}$ (recall Definition 3.1), the set of split vertices corresponding to the inter-cluster edges
@@
    $\alpha$-expands in $G[S]^{\prime}$ for $\alpha=\Omega\!\left(\frac{1}{\log n}\right)$.
@@
Note that Theorem 4.1 has $\alpha=\Omega\!\left(\frac{1}{\log n}\right)$, instead of $\alpha=\Omega\!\left(\frac{1}{\log^{2}n}\right)$ in [27, Theorem 3.1], which leads (as shown below) to the improvement by a factor of $\log n$ in the quality of the tree cut-sparsifier. The partitioning algorithm in Theorem 4.1 is independent of the edges outside $G[S]$. Therefore, we can restrict our attention to the induced subgraph $G[S]$: in the rest of this section, $G=(V,E)$ refers to $G[S]$, and $n,m$ denote the number of vertices and edges, respectively, in $G[S]$. Following the notation of [27], we say that a subset $F\subseteq E$ of edges induces a balanced clustering in $G$, if deleting $F$ from $G$ leaves connected components with at most $\frac{2}{3}|V|$ vertices in each.
To establish Theorem 4.1, we state the following lemma, which improves and unifies Lemmas 3.1 and 3.2 in [27]. We defer the proof of Lemma 4.2 to the end of the section.
Lemma 4.2 (Improved version of [27, Lemmas 3.1 and 3.2]).
@@
    an edge set $\tilde{F}=\tilde{A}\cup\tilde{C}$, such that $\tilde{F}$ induces a balanced clustering, and with high probability $X_{\tilde{A}}$, $\Omega(1/\log n)$-expands in $G^{\prime}$. Additionally there exists a flow in $G^{\prime}$ from nodes in $X_{\tilde{C}}$ to $X_{\tilde{A}}$ that can be routed with congestion $O(1)$ and:
@@
The following lemma shows that in Case (2) of Lemma 4.2, the set $X_{\tilde{F}}$, $\Omega(1/\log n)$-expands in $G^{\prime}$.
Lemma 4.3.
Let $G$ be a graph, and let $A,C\subseteq V$. Assume that AA $\phi$-expands in $G$, and additionally that there exists a flow routable with congestion $c$, such that each vertex $v\in C$ sends $1$ unit of flow, and each vertex $u\in A$ receives at most $a$ units (other vertices receive no flow). Then $A\cup C$ $\frac{\phi}{2\cdot(2+2a+c\phi)}$-expands in $G$.
Before proving Lemma 4.3, we show how Lemmas 4.2 and 4.3 imply Theorem 4.1.
@@
Once Lemma 4.2 terminates with Case (2), $X_{\tilde{A}}$, $\Omega(1/\log n)$-expands in $G^{\prime}$, and we can route in $G^{\prime}$ with congestion $O(1)$ a flow such that each node in $X_{\tilde{C}}$ sends $1$ unit of flow and each node in $X_{\tilde{A}}$ receives at most $1$ unit of flow. By Lemma 4.3, this implies that edges in $\tilde{F}=\tilde{A}\cup\tilde{C}$ satisfy that $X_{\tilde{F}}$, $\Omega(1/\log n)$-expands in $G^{\prime}$. Additionally, $\tilde{F}$ induces a balanced clustering whose connected components will be $\{Z_{1},\ldots,Z_{z}\}$. [^8] After computing $\tilde{F}$ and the resulting partition $\{Z_{1},\ldots,Z_{z}\}$, we re-define $F$ to be the new set of inter-cluster edges $F\mathrel{\mathop{:}}=\{(u,v)\mid u\in Z_{i},v\in Z_{j},(u,v)\in E,i\neq j\}$. This set satisfies $F\subseteq\tilde{F}$ and thus $\Omega(1/\log n)$-expands in $G[S]^{\prime}$ (see Fact 3.8). This proves Theorem 4.1. ∎
Next, we prove the deferred lemmas, starting with Lemma 4.3 and concluding with Lemma 4.2. We begin by giving intuition for Lemma 4.3. By Remark 3.6, it is enough to prove that $G$ $\frac{\phi}{2\cdot(2+2a+c\phi)}$-respects the all-to-all flow problem between the nodes of $A\cup C$ (see Definition 2.2).
Using the fact that $A$ expands in $G$ (thus $G$ respects the all-to-all flow problem on $A$) and that there is a low-congestion flow $f$ from $C$ to $A$ (in particular $G$ respects the flow problem that moves mass from $C$ to $A$ according to $f$), we observe that the all-to-all demand matrix on $A\cup C$ can be realized by the following sequence of demand transitions, each of which is respected by $G$:
@@
    First, “move” the demands in $C$ to $A$ according to $f$.
@@
    Then, spread the demands in $A$ uniformly (using the fact that $A$ expands in $G$).
@@
    Finally, “move” part of the (uniform) demands in $A$ back to $C$ according $f$ (in reverse).
We now formalize this argument.
Proof of Lemma 4.3.
Consider the all-to-all demand matrix $Q_{A\cup C}$ on $A\cup C$ in $G$ (recall Definition 2.2). By Remark 3.6, it suffices to show that $G$ respects $Q_{A\cup C}$. Let $P_{1}\mathrel{\mathop{:}}=P_{A\cup C}$ be the corresponding valid demand state (see Remark 3.14). Note that for each $u\in A\cup C$, $\left\|{P_{1}}(u)\right\|_{1}=2\cdot\left(1-\frac{1}{|A\cup C|}\right)\leq 2$.
@@
We use the given flow from $C$ to $A$, which corresponds to a demand matrix denoted by $Q$, to transport the demand state vectors of the vertices in $C\setminus A$ to vertices in $A$: We scale the demands of $Q$ such that each $u\in C\setminus A$ sends $2\cdot\left(1-\frac{1}{|A\cup C|}\right)$ units and each $u\in C\cap A$ sends 0. That is, we multiply the rows of $Q$ corresponding to $C\setminus A$ by $2\cdot\left(1-\frac{1}{|A\cup C|}\right)$ and rows of $C\cap A$ are zeroed out. Let $Q^{\prime}$ be the scaled demand matrix. Because each of the rows was scaled by at most $2$, $Q^{\prime}$ can be routed with congestion $2c$ and every vertex $v\in A$ receives at most $2a$ units. Next, consider the updated demand state $P_{2}\mathrel{\mathop{:}}=P_{1}^{\uparrow Q^{\prime}}$ (recall Definition 3.17). By Lemma 3.20(2), $P_{2}$ is supported only on $A$. Moreover, for every $u\in A$, the load satisfies $\left\|{P_{2}}(u)\right\|_{1}\leq 2+2a$. Indeed, by triangle inequality,
$\left\|P_{2}(u)\right\|_{1}$ $\leq\left\|P_{1}(u)-\frac{P_{1}(u)}{\left\|P_{1}(u)\right\|_{1}}\cdot\sum_{v\in V}{Q^{\prime}(u,v)}\right\|_{1}+\left\|\sum_{v\in V}{\frac{P_{1}(v)}{\left\|P_{1}(v)\right\|_{1}}\cdot Q^{\prime}(v,u)}\right\|_{1}$
$\leq\left\|P_{1}(u)\right\|_{1}+\sum_{v\in V}Q^{\prime}(v,u)\leq\left\|P_{1}(u)\right\|_{1}+2a\leq 2+2a.$
Because $Q^{\prime}$ can be routed with congestion $2c$, it follows (see Fact 2.7) that $G$, $\frac{1}{2c}$-respects $Q^{\prime}$. Thus, by Fact 3.19(2), for any cut $(U,W=V\setminus U)$ we have $\operatorname{dem}_{(P_{1}-P_{2})}(U,W)\leq 2c\cdot\operatorname{cap}_{G}(U,W)$.
@@
Next, since AA $\phi$-expands in $G$, Remark 3.6 implies that $G$ $\frac{\phi}{2}$-respects the all-to-all flow problem on $A$, whose demand matrix is denoted by $Q_{A}$. By scaling the rows of $Q_{A}$ by at most $2+2a$, we can get a demand matrix which routes $\left\|{P_{2}}(u)\right\|_{1}$ units out of each $u\in A$, and spreads them uniformly on $A$. This scaled problem is $\frac{\phi}{4+4a}$-respected by $G$. Updating $P_{2}$ using this scaled flow gives $P_{3}$, which by Lemma 3.20, has identical demand vectors for all $u\in A$. Because $P_{1}$ is a valid demand state, so are $P_{2}$ and $P_{3}$ (see Fact 3.19(1)). Since all non-zero demand vectors in $P_{3}$ are equal, this means that $P_{3}\equiv 0$.
@@
Similar to before, we get that for any cut $(U,W)$ we have $\operatorname{dem}_{(P_{2}-P_{3})}(U,W)\leq\frac{4+4a}{\phi}\cdot\operatorname{cap}_{G}(U,W)$. Therefore, using Fact 3.16, for any cut $(U,W)$, we have
$\operatorname{dem}_{Q_{A\cup C}}(U,W)$ $=\operatorname{dem}_{P_{1}}(U,W)\leq\operatorname{dem}_{(P_{1}-P_{2})}(U,W)+\operatorname{dem}_{(P_{2}-P_{3})}(U,W)$
$\leq\left(2c+\frac{4+4a}{\phi}\right)\cdot\operatorname{cap}_{G}(U,W)=\frac{2\cdot(2+2a+c\phi)}{\phi}\cdot\operatorname{cap}_{G}(U,W),$
@@
Figure 2: Partitioning of cluster $S$. Figures (a)-(c) represent $G$, figures (d)-(f) represent the split graph. Black ovals depict vertices and squares depict split nodes. Figures $(b)$ and $(e)$ correspond to the first partition. Blue edges (and blue squares) depict inter-cluster edges of the first partition (that is, $F$ and $X_{F}$, respectively). Figure $(f)$ depicts the second partition (which is computed in $G^{\prime}[S^{\prime}]$) (recall $S^{\prime}=S\cup\{x_{e}\mid e\cap S\geq 1\}$). The core $L$ is enveloped in gold. Split nodes in $X_{Y}$ are filled with gold. Split nodes in $X_{F\cap Y}$ are colored in blue and gold.
To prove Lemma 4.2 we need the following claim. It is a slightly more general rephrasing of Claim 1 of [27]. We provide its proof for completeness.
Claim 4.4 ([27, Claim 1]).
@@
Given the subset $F\subseteq E$ that induces a balanced clustering, we use Theorem 3.9 on $G^{\prime}$ with $\phi\mathrel{\mathop{:}}=\Theta\left(\frac{1}{\log n}\right)$ and $\mu\mathrel{\mathop{:}}=\mu_{X_{F}}$. It terminates with one of three cases:
@@
    for some constant $c_{1}>0$. Let $C^{\prime}\subseteq E^{\prime}$ be the set of edges that cross the cut $(A^{\prime},\bar{A^{\prime}})$ in $G^{\prime}$. Let $C\subseteq E$ be the corresponding set in $G$: $C\mathrel{\mathop{:}}=\{e\in E\mid\exists c\in C^{\prime}:x_{e}\in c\}$. We have $|C|\leq|C^{\prime}|$. Inequality (1) is equivalent to $|C^{\prime}|\leq\frac{1}{2}\min(|X_{F}\cap A^{\prime}|,|X_{F}\setminus A^{\prime}|)$. Let $A\subseteq E$ be the set of edges corresponding to $A^{\prime}$: $A\mathrel{\mathop{:}}=\{e\in E\mid x_{e}\in A^{\prime}\}$. We have $|F\cap A|=|X_{F}\cap A^{\prime}|$ and $|F\setminus A|=|X_{F}\setminus A^{\prime}|$. Therefore,
@@
    We find a cut $(A^{\prime},\bar{A^{\prime}})$ in $G^{\prime}$ with $0<|X_{F}\setminus A^{\prime}|\leq\frac{|F|}{2}$, $\Phi_{G^{\prime}}^{X_{F}}(A^{\prime},\bar{A^{\prime}})=O(\phi\log n)$, and with high probability, $A^{\prime}$ $\Omega(\phi)$-expands with respect to $X_{F}$ in $G$. If $|X_{F}\setminus A^{\prime}|>\frac{|F|}{\log n}$, then we complete the proof as in Case (2). Therefore, assume that $0<|X_{F}\setminus A^{\prime}|\leq\frac{|F|}{\log n}$. As $\phi=\Theta\left(\frac{1}{\log n}\right)$, denote by $c_{2}>0$ the constant such that $\Phi_{G^{\prime}}^{X_{F}}(A^{\prime},\bar{A^{\prime}})\leq c_{2}$. Like in the previous case, denote by $C^{\prime}\subseteq E^{\prime}$ the set of edges that cross the cut $(A^{\prime},\bar{A^{\prime}})$ in $G^{\prime}$. Let $C\subseteq E$ be the corresponding set in $G$: $C\mathrel{\mathop{:}}=\{e\in E\mid\exists c\in C^{\prime}:x_{e}\in c\}$. We have $|C|\leq|C^{\prime}|$. Because the cut is sparse with respect to $\mu_{X_{F}}$, $|C^{\prime}|\leq c_{2}\cdot|X_{F}\setminus A^{\prime}|$.
    From Theorem 3.10, we get that there exists a flow in $G^{\prime}$, routable with congestion $O(1)$, such that each split-node in $X_{C}$ sends $1$ unit of flow and each node $v\in A^{\prime}$ receives at most $O(\phi\cdot\mu(v))$ units. Recall that $\phi=\Theta\left(\frac{1}{\log n}\right)$, so for large enough $n$, we may assume that each node receives at most $\mu(v)$ units of flow. In particular, only split nodes $v\in X_{F}$ can receive flow, and each receives at most one unit.
@@
      Assume $\tilde{R}\cup C$ induces a balanced clustering. In this case, we will set $\tilde{F}\mathrel{\mathop{:}}=\tilde{R}\cup C$, and claim that we get Case (1) of the lemma. Indeed, we have $|\tilde{R}\cup C|\leq|\tilde{R}|+|C|=|F\setminus A|+|C|\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu1\mkern-1.5mu)}}{{\leq}}(c_{2}+1)\cdot|X_{F}\setminus A^{\prime}|\leq(c_{2}+1)\cdot\frac{|F|}{\log n}<|F|\cdot(1-\frac{1}{\log n})$, where Inequality $(1)$ holds since $|C|\leq|C^{\prime}|\leq c_{2}\cdot|X_{F}\setminus A^{\prime}|$, and the last inequality holds for large enough $n$.